# A Power Method for Computing Singular Value Decomposition

### v2

@@ 3.1 Principal component analysis
In the SVD analysis, no assumption was used on the statistical properties for the entries of the matrix $\mathbf{X}$. In practical applications, the row $\mathbf{x}_{i}$, $i=1,2,\ldots,m$, of this matrix is a measurement sample of a phenomenon. Let us assume that the $n$-dimensional vector $\mathbf{x}$ is an observation of a random vector $\mathcal{X}$ which entries are the variables of the phenemenon under study. The mean and the covariance of this random vector are: $\mathbb{E}[\mathcal{X}]\triangleq\boldsymbol{\mu}$ and $\mathbb{V}[\mathcal{X}]=\mathbb{E}[(\mathcal{X}-\boldsymbol{\mu})(\mathcal{X}-\boldsymbol{\mu})^{T}]\triangleq\boldsymbol{\Sigma}$. The diagonal entries of the matrix $\boldsymbol{\Sigma}$ are the variances of the $\mathcal{X}$ entries: $cov(\mathbf{x}_{j},\mathbf{x}_{j})=\sigma_{j}^{2}$, $j=1,2,\ldots,n$. The principal component analysis (PCA) consists of the eigendecomposition of the covariance matrix, followed by a representation of the observations and variables in the reduced space axes, [19, 20]. Very often, the random vector $\mathcal{X}$ is assumed to be of zero mean. When the entries of $\mathcal{X}$ are heterogenous, there can be a large gap between the amplitudes of the variances $\sigma_{j}^{2}$, $j=1,2,\ldots,n$. In this case the role of the variables with small variances is hidden in the interpretation of the PCA results. To have homogeneous variance data, the centered value of each entry should be divided by the corresponding standard deviation $\sigma_{j}$. The covariance matrix becomes a correlation matrix which has only ones as diagonal entries.
3.1.1 PCA for finite number of observations
When $m$ observations or objects are available, only estimates can be obtained for the mean vector $\boldsymbol{\mu}$ and the covariance matrix $\boldsymbol{\Sigma}$. The sample mean $\bar{\mathbf{x}}$ and the sample variance $s_{j}^{2}$ of the $j$-th entry of the observations $\mathbf{x}_{i}$, $i=1,2,\ldots,m$, are estimated as:
@@ 3.1.1 PCA for finite number of observations
From the entries $x_{ij}$ of the matrix $\mathbf{X}$, the entries $x_{ij}^{c}$ of the centered matrix $\mathbf{X}_{c}$ and the entries $x_{ij}^{s}$ of the standardized matrix $\mathbf{X}_{s}$ are obtained as ($i=1,2,\ldots,m$, $j=1,2,\ldots,n$):
$\mathbf{X}_{c}$ ::: $x^{c}_{ij}=x_{ij}-\bar{x}_{j}$
$\mathbf{X}_{s}$ ::: $x^{s}_{ij}=(x_{ij}-\bar{x}_{j})/s_{j}$
Estimates for the sample covariance and the sample variance matrices are obtained using $\mathbf{X}_{c}^{\mathsf{T}}\mathbf{X}_{c}/m$ and $\mathbf{X}_{s}^{\mathsf{T}}\mathbf{X}_{s}/m$, respectively. Using the expressions between the SVD factors and those of the eigendecomposition, the PCA result can be obtained from the SVD results of the matrix $\mathbf{X}_{c}$ or the matrix $\mathbf{X}_{s}$.
3.1.2 Representation of objects or variables in reduced space
The matrices $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ and $\mathbf{XX}^{\mathsf{T}}$ have the same eigenvalues and related eigenvectors. The eigenvalue $\lambda$ allows to write:
@@ 3.1.2 Representation of objects or variables in reduced space
where $\mathbf{v}$ and $\mathbf{u}$ are the eigenvectors. Relation (25) allows to write: $\mathbf{XX}^{\mathsf{T}}\mathbf{Xv}=\lambda\mathbf{Xv}$, meaning that $\mathbf{u}=\mathbf{Xv}$. Relation (26) can be used in a similar way to show that $\mathbf{v}=\mathbf{X}^{\mathsf{T}}\mathbf{u}$. Let $r$ be the reduced dimension used in the PCA, typically $r=2$ or $3$ for a visual representation. From the PCA results of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$, expressions for the objects in the $r$-dimension variable space are $\mathbf{X}\mathbf{V}_{r}$, where $\mathbf{V}_{r}$ is a matrix with the $r$ largest eigenvectors. Using the expression for the matrix $\mathbf{U}$ in relation (14), expressions for the variables in the $m$-dimensional space are obtained as: $\mathbf{X}^{\mathsf{T}}\mathbf{XV}_{r}\boldsymbol{\Lambda}_{r}^{-1/2}$, where $\boldsymbol{\Lambda}_{r}^{-1/2}$ is the inverse of the diagonal matrix obtained using the $r$ largest nonzero eigenvalues.
The PCA results plot for the iris data is shown in Figure 2.
Figure 2: Plot of the iris observations in the first and third principal components axes.
@@ 3.2 Autoencoder
An autoencoder (AE) is a neural network (NN) model which allows to encode input, store the result internally, and finally decode the stored result to obtain an output. The output is not necessary strictly equal to the input. In comparison, the SVD leads to the matrices $\mathbf{U}$, $\mathbf{D}$ and $\mathbf{V}$ which are an internal representation of the input matrix $\mathbf{X}$. Then, using a rank $r$ value, an approximation of the input is obtained: $\hat{\mathbf{X}}=\mathbf{U}_{r}\mathbf{D}_{r}\mathbf{W}_{r}^{\mathsf{T}}$. This approximation matrix is of rank $r$ and the SVD uses linear operations to solve the dimensionality reduction problem or data compression. The data can also be compressed using a feedforward NN model with few layers, [30, 4, 22]. A NN layer consists of a weighted combination of an input, $\mathbf{x}$, followed by a transformation function $\varphi(.)$ to obtain an output: $\mathbf{u}=\mathbf{x}^{\mathsf{T}}\mathbf{W}+\mathbf{w}_{o}$; $\mathbf{v}=\varphi(\mathbf{u})$, where $\mathbf{W}$ is weight matrix and $\mathbf{w}_{o}$ is a threshold or bias vector. The transformation operates on each entry of the vector $\mathbf{u}$. Three examples of the transformation function are the sigmoid, $\sigma(u)=1/(1+e^{-u})$, the rectified linear unit, $ReLU(u)=max(0,u)$, and the identity $I(u)=u$. A two layers feedforward NN model has been used in [22] to perform a nonlinear SVD for obtaining an approximation of the input $\mathbf{X}$. Each row of the matrix $\mathbf{X}$ is an $n$-dimensional vector $\mathbf{x}_{i}$, $i=1,2,\ldots,m$, and the output of the encoding layer is obtained as:
@@
where $\mathbf{W}_{e}$ is a weight matrix of size $n\times r$ and $\mathbf{w}_{e}$ is a bias vector of order $n$. The output of the decoding layer is obtained as:
$\hat{\mathbf{x}}_{i}=\sigma\left(\mathbf{v}_{i}^{\mathsf{T}}\mathbf{W}_{d}+\mathbf{w}_{d}\right)=\sigma\left(\sigma\left(\mathbf{x}_{i}^{\mathsf{T}}\mathbf{W}_{e}+\mathbf{w}_{e}\right)^{\mathsf{T}}\mathbf{W}_{d}+\mathbf{w}_{d}\right)$
where $\mathbf{W}_{d}$ is a weight matrix of size $r\times n$ and $\mathbf{w}_{d}$ is a bias vector of order $n$. The input $\mathbf{x}_{i}$ and its approximation by the decoder output are used to define an $\ell_{2}$ norm:
@@
From the $m$ observations, the weight matrices, $\mathbf{W}_{e}$ and $\mathbf{W}_{d}$, and the bias vectors, $\mathbf{w}_{e}$ and $\mathbf{w}_{d}$ are obtained after minimization of the $\ell_{2}$-norm using a gradient search optimization algorithm. Instead of using a single layer in the encoder, $k$ layers, $k>1$, can be used to progressively increase the compression rate. In this case, an inverse operation is done using $k$ decoding layers. The approximation matrix obtained using the feedforward NN model is not necessary of reduced rank even the number of columns of the encoder weight matrix is $r<n$. However, equivalent approximation to SVD result can be obtained using the AE model. The difference matrix $\mathbf{X}-\hat{\mathbf{X}}$ can be viewed as the noise part of the data. Therefore, the approximation $\hat{\mathbf{X}}$ can be used in subsequent supervised or unsupervised data analysis methods.
3.2.1 Identity transformation function
The identity transformation function has been used in [33, 36]. The transpose of the matrix $\mathbf{X}$ is used in [33] leading to $\|\mathbf{X}-\mathbf{WW}^{\mathsf{T}}\mathbf{X}\|$ while the constrained optimization objective function of the relation (16) is used in [36, Chapter 7]. In these two references, the classical gradient search optimization method is used to obtain the matrix $\mathbf{W}$. The algorithm 1 solves the same problem by converting the gradient search into a power method algorithm and adds a convergence acceleration through the choice of the parameter $q$. When an estimation is available for the weight matrix $\mathbf{W}$, then an proximation of the input matrix $\mathbf{X}$ is: $\hat{\mathbf{X}}=\mathbf{XWW}^{\mathsf{T}}$. Therefore, an approximation of the input vector $\mathbf{x}_{i}$ is: $\hat{\mathbf{x}}_{i}=\mathbf{x}_{i}^{\mathsf{T}}\mathbf{WW}^{\mathsf{T}}$.
MNIST data.
For illustration, the MNIST data, [26], was used. This dataset consists of $60,000$ training images and $10,000$ test images. Each image of size $28\times 28$ is an observation or vector of dimension $784$. The entries of this vector are pixels which values are from the set $\{0,1,\ldots,255\}$. The behavior of the singular values and the cumulative sum of the reconstruction percentage rate of the training images data are show in Figure 3. The last $72$ singular values are zero. The total information captured by the first $50$, $100$ or $256$ largest singular values are $45.41\%$, $60.45\%$ and $82.63\%$, respectively.
@@ 3.2.1 Identity transformation function
Figure 3: Log scale singular values and cumulative percentage reconstruction rate
The largest $256$ singular values were computed using the training images data. The weight matrix obtained was used to estimate approximates for the first $5$ test images. The results are shown in Figure 4, where the first column is the test image and the $3$ following columns are for the approximation using the first $50$, $100$ or $256$ singular values, respectively. The compression rates for these approximations are $15.68$, $7.84$ and $3.06$, respectively.
Figure 4: First 5 test images, from left to right: input, use of the $50$, $100$ or $256$ largest singular values to obtain approximation.
@@ 4 Conclusion
A new method is proposed for computing the singular values of a matrix. This method allows to compute all or a subset of the largest singular values and associated vectors. It can be applied to small or high dimensional dataset. The accuracy of the results depends on the user tolerance setting. More accuracy may necessitate a long runtime depending on the behavior of the singular values decay. The proposed method especially consists of searching for an eigendecomposition of a square symmetric matrix by using a gradient method to minimize a Frobenius norm. The method was applied to principal component analysis and to neural network autoencoder model.
@@
The proposed method is implemented in the psvd R package available on comprehensive R archive network [35]: http://cran.r-project.org/web/packages/psvd/. In the first version, the power parameter $q$ is fixed to $2$ and only a sample from standard normal distribution is used as initial projection matrix. The basic power method algorithm is given below to show the similarity with the proposed algorithm. For completeness, a numerically stable Gram-Schmidt orthogonalization algorithm is provided below. The iterative part of the proposed algorithm is the more runtime consuming. A C code is used for this part in the psvd R package. Reader can use her/his preferred computing language to implement all these algorithms.
@@
Result: orthonormal vector $\mathbf{v}$

### v4

@@ 4.1 Principal component analysis
In the SVD, no assumption is used on the statistical properties for the entries of the matrix $\mathbf{X}$. In practical applications, the entries of the row $\mathbf{x}_{i}$, $i=1,2,\ldots,m$, are the measurements associated with $n$ attributes or features. Let us assume that the $n$-dimensional vector $\mathbf{x}$ is a sample of a random vector $\mathcal{X}$. The mean and the covariance of this random vector are: $\mathbb{E}[\mathcal{X}]\triangleq\boldsymbol{\mu}$ and $\mathbb{V}[\mathcal{X}]=\mathbb{E}[(\mathcal{X}-\boldsymbol{\mu})(\mathcal{X}-\boldsymbol{\mu})^{T}]\triangleq\boldsymbol{\Sigma}$. The diagonal entries of the matrix $\boldsymbol{\Sigma}$ are the attributes variances: $cov(\mathbf{x}_{j},\mathbf{x}_{j})=\sigma_{j}^{2}$, $j=1,2,\ldots,n$. The principal component analysis (PCA) consists of the EVD of the covariance matrix, [30, 31]. When the $\mathcal{X}$’s entries are heterogenous, there can be a large gap between the amplitudes of the variances $\sigma_{j}^{2}$, $j=1,2,\ldots,n$, leading a misinterpretation of the PCA result. To have homogeneous variances, the samples are standardized and the correlation matrix is used in the PCA.
PCA for finite number of samples
When $m$ samples are available, the sample mean $\bar{\mathbf{x}}$ and the sample variance $s_{j}^{2}$ of the $j$-th entry of the sample $\mathbf{x}_{i}$, $i=1,2,\ldots,m$, are estimated:
@@
Then, the entries $x_{ij}^{c}$ of the centered matrix $\mathbf{X}_{c}$ and the entries $x_{ij}^{s}$ of the standardized matrix $\mathbf{X}_{s}$ are obtained as ($i=1,2,\ldots,m$, $j=1,2,\ldots,n$):
$\mathbf{X}_{c}$ $:$ $x^{c}_{ij}=x_{ij}-\bar{x}_{j}$
$\mathbf{X}_{s}$ $:$ $x^{s}_{ij}=(x_{ij}-\bar{x}_{j})/s_{j}$
Estimates for the sample covariance and the sample variance matrices are obtained as $\mathbf{X}_{c}^{\mathsf{T}}\mathbf{X}_{c}/m$ and $\mathbf{X}_{s}^{\mathsf{T}}\mathbf{X}_{s}/m$, respectively. Using the expressions between the SVD factors and those of the EVD, the PCA result can be obtained by using the matrix $\mathbf{X}_{c}$ or the matrix $\mathbf{X}_{s}$.
Representation of the samples or the attributes in a reduced space
The matrices $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ and $\mathbf{XX}^{\mathsf{T}}$ have the same eigenvalues and related eigenvectors:
@@
where $\mathbf{v}$ and $\mathbf{u}$ are the eigenvectors. Relation (26) allows to write: $\mathbf{XX}^{\mathsf{T}}\mathbf{Xv}=\lambda\mathbf{Xv}$, meaning that $\mathbf{u}=\mathbf{Xv}$. Relation (27) can be used in a similar way to show that $\mathbf{v}=\mathbf{X}^{\mathsf{T}}\mathbf{u}$. Let $r$ be the reduced dimension used in the PCA, typically $r=2$ or $3$ for a visual representation. From the PCA result of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$, expressions for the samples in the $r$-dimension attribute space are $\mathbf{X}\mathbf{V}_{r}$, where $\mathbf{V}_{r}$ is a matrix with the $r$ first largest eigenvectors. Using the expression for the matrix $\mathbf{U}$ in the relation (6), the attributes are expressed in the $m$-dimensional space as: $\mathbf{X}^{\mathsf{T}}\mathbf{XV}_{r}\boldsymbol{\Lambda}_{r}^{-1/2}$, where $\boldsymbol{\Lambda}_{r}^{-1/2}$ is the square root inverse of the diagonal matrix associated with the $r$ first largest nonzero eigenvalues. The PCA result plot for the iris data is shown in Figure 2 where the $r=3$ first singular values and singular vectors were computed, the default settings were used for the other parameters.
Figure 2: Plot of the iris samples in the first and third principal components axes.
@@ 4.2 Autoencoder
An autoencoder is a neural network (NN) model which allows to encode input, store the result internally, and finally decode the stored result to obtain an output. The output is not necessary strictly equal to the input. In comparison, the SVD leads to the matrices $\mathbf{U}$, $\mathbf{D}$ and $\mathbf{V}$ which are an internal representation of the input matrix $\mathbf{X}$. Then, using the $r$ first largest singular values, a low-rank approximation is obtained: $\hat{\mathbf{X}}=\mathbf{U}_{r}\mathbf{D}_{r}\mathbf{W}_{r}^{\mathsf{T}}$. Thus, the SVD uses linear algebra operations to solve the dimensionality reduction problem or data compression. The data can also be compressed using a feedforward NN model with few layers, [49, 8, 35]. A NN layer is a weighted combination of an input, $\mathbf{x}$, followed by a transformation function $\varphi(.)$ to obtain an output: $\mathbf{u}=\mathbf{x}^{\mathsf{T}}\mathbf{W}+\mathbf{w}_{o}$; $\mathbf{v}=\varphi(\mathbf{u})$, where $\mathbf{W}$ is a weight matrix and $\mathbf{w}_{o}$ is a threshold or bias vector. The transformation operates on each entry $u$ of the vector $\mathbf{u}$. Three examples of the transformation function are the sigmoid, $\sigma(u)=1/(1+e^{-u})$, the rectified linear unit, $ReLU(u)=max(0,u)$, and the identity $I(u)=u$. A two layers feedforward NN model has been used in [35] to perform a nonlinear SVD for an approximation of $\mathbf{X}$. The encoder output $\mathbf{v}_{i}$ is obtained for each input $\mathbf{x}_{i}$, $i=1,2,\ldots,m$, as:
@@
where $\mathbf{W}_{e}$ is the encoder weight matrix of size $n\times r$ and $\mathbf{w}_{e}$ is a bias vector of order $r$. The output of the decoding layer is obtained as:
$\hat{\mathbf{x}}_{i}=\sigma\left(\mathbf{v}_{i}^{\mathsf{T}}\mathbf{W}_{n}+\mathbf{w}_{n}\right)=\sigma\left(\sigma\left(\mathbf{x}_{i}^{\mathsf{T}}\mathbf{W}_{e}+\mathbf{w}_{e}\right)^{\mathsf{T}}\mathbf{W}_{n}+\mathbf{w}_{n}\right)$
where $\mathbf{W}_{n}$ is a weight matrix of size $r\times n$ and $\mathbf{w}_{n}$ is a bias vector of order $n$. The input $\mathbf{x}_{i}$ and its approximation, the decoder output, can be used to define a $\ell_{2}$-norm:
@@
From $m$ samples, the weight matrices, $\mathbf{W}_{e}$ and $\mathbf{W}_{n}$, and the bias vectors, $\mathbf{w}_{e}$ and $\mathbf{w}_{n}$ are obtained after minimization of the $\ell_{2}$-norm using a gradient search optimization algorithm. The approximation matrix obtained using a nonlinear activation feedforward NN model is not necessary of reduced rank even the number of columns of the encoder weight matrix is $r<n$. However, equivalent approximation to SVD result can be obtained using the autoencoder model. The difference matrix $\mathbf{X}-\hat{\mathbf{X}}$ can be viewed as the noise part of the data and the approximation $\hat{\mathbf{X}}$ can be used in subsequent supervised or unsupervised data analysis methods.
Identity transformation function
The identity transformation function has been used in [52, 55]. The transpose of the matrix $\mathbf{X}$ is used in [52] leading to $\|\mathbf{X}-\mathbf{WW}^{\mathsf{T}}\mathbf{X}\|$ while the constrained optimization objective function of the relation (16) is used in [55, Chapter 7]. In these references, the classical gradient search optimization method is used to obtain the matrix $\mathbf{W}$. The $k$ columns of this matrix are not necessary the $k$ first singular vectors of the data matrix. For the proposed method, the $k$ first largest singular vectors are the $k$ columns of the autoencoder model weight matrix. When an estimation is available for $\mathbf{W}$, then approximations for the matrix $\mathbf{X}$ and the vector $\mathbf{x}_{i}$ are $\hat{\mathbf{X}}=\mathbf{X}\hat{\mathbf{W}}\hat{\mathbf{W}}^{\mathsf{T}}$ and $\hat{\mathbf{x}}_{i}=\mathbf{x}_{i}^{\mathsf{T}}\hat{\mathbf{W}}\hat{\mathbf{W}}^{\mathsf{T}}$, respectively.
MNIST data
This dataset consists of $60,000$ training images and $10,000$ test images, [40]. Each image of size $28\times 28$ is a sample, a vector of order $784$. The entries of this vector are pixels which values are from the set $\{0,1,\ldots,255\}$. The behavior of the singular values and the cumulative sum of the reconstruction percentage rate of the training images data are show in Figure 3. The last $72$ singular values are zero. The total information captured by the $50$, $100$ or $256$ first largest singular values are $45.41\%$, $60.45\%$ and $82.63\%$, respectively.
@@
Figure 3: Log scale singular values and cumulative percentage reconstruction rate for the MNIST images
The $256$ first largest singular values were computed using the training images data and each method. The results are shown in Table 5. In the first part of this Table, the default parameter itmax (200) is used while itmax = 150(a), 100(b) and 50(c), are used for the 3 last rows results (proposed method only). Like for the trachea dataset, the rsvd method has the highest MSE value, with a reconstruction rate almost similar to that of the other methods. When accurate singular values are not required, a fast runtime is obtained by using a small value for the parameter itmax in the algorithm 1.
Table 5: Calculation of the $256$ first singular values for the MNIST images data using 6 methods: runtime, reconstruction rate and MSE comparison results
method runtime reconst.rate MSE
svd 85.26 82.63 -
propack.svd 71.35 82.63 <<1e-10
RSpectra 47.50 82.63 <<1e-10
irlba 219.74 82.63 5.99e-6
rsvd 87.68 82.35 16,631.1
psvd 82.75 82.63 29.16
psvd(a) 69.72 82.63 71.43
psvd(b) 57.71 82.63 173.8
psvd(c) 45.35 82.63 2,777.4
The weight matrix obtained using the proposed method was used to have the $5$ first test images approximations. The results are shown in Figure 4, where the first column is the test image and the $3$ following columns are the results using the $50$, $100$ or $256$ first largest singular values approximations, respectively. The compression rates for these approximations are $15.68$, $7.84$ and $3.06$, respectively.
Figure 4: 55 first test images, from left to right: input, use of the $50$, $100$ or $256$ first largest singular values to obtain approximation.
@@ 5 Conclusion
A new method is proposed for computing the singular values of a real-valued entries matrix. This method consists of using a gradient search optimizer to minimize a Frobenius norm and allows to compute all or a subset of the $k$ first largest singular values and singular vectors. The accuracy of the results depends on the user settings. More accuracy may necessitate a long runtime depending on the behavior of the singular values decay. A parallel implementation may allow to speed up the runtime of the proposed SVD calculation algorithm. The method was applied to principal component analysis and to neural network autoencoder model. An independent optimization method is often used in the literature to get the singular vectors which are not necessary the $k$ first columns of the encoder weight matrix unlike for the method proposed.
@@
The proposed method is implemented in the psvd R package available on comprehensive R archive network [54]: http://cran.r-project.org/web/packages/psvd/. The power parameter $q$ can be chosen in the set $\{2,3,4\}$, a sample from standard normal distribution is used as initial projection matrix. The basic power method algorithm is given below to show the similarity with the proposed algorithm. For completeness, a numerically stable Gram-Schmidt orthogonalization algorithm is provided below. The iterative part of the proposed algorithm is the more runtime consuming. A C code is used for this part in the psvd R package. Reader can use her/his preferred computing environment to implement these algorithms.
@@
Result: orthogonal vector $\mathbf{v}$